OpenStudy (anonymous):

Paula ran for 195 minutes this weekend. She ran 25 minutes more on Saturday than Sunday. How long did she run each day?

OpenStudy (anonymous):

x + x + 25 = 195 x = 170/2 = 85 x + 25 = 110 she ran 110 minutes on S aturday and 85 mins Sunday